# Fibered ribbon pretzels

### v1

@@ Abstract.
We classify fibered ribbon pretzel knots up to mutation. The classification is complete, except perhaps for members of Lecuona’s “exceptional” family of [Lec15]. The result is obtained by combining lattice embedding techniques with Gabai’s classification of fibered pretzel knots, and exhibiting ribbon disks, some of which lie outside of known patterns for standard pretzel projections.
@@ 1. Introduction
The intersection of fiberedness and ribbon/sliceness is of particular interest, in large part due to work of Casson and Gordon [CG83], who observe that existence of a ribbon disc implies a surjection at the level of $\pi_{1}$ from the knot complement to the disk complement, which of course by definition implies sliceness. Labelling this algebraic condition homotopically ribbon, we have thus the following inclusions:
$\{\mathrm{ribbon}\}\subseteq\{\mathrm{homotopically\ ribbon}\}\subseteq\{\mathrm{slice}\}$
Moreover, in the case of fibered knots, Casson and Gordon show that homotopical ribonness admits a simple characterisation in terms of the fibration, namely that its monodromy extends over a handlebody. Investigation of the details of this result led Miller [Mil21, Remark 6.4] to study fibered ribbon 2-bridge knots, concluding that there are exactly 5 of these.
@@
To state our results, some conventions, standard when discussing pretzel knots, need to be established. Firstly, as we describe in detail in the next section, such knots can be encoded by an ordered tuple of integers. Mutations of the knot correspond to re-orderings of these parameters; as mutation preserves most known obstructions to sliceness, we will only consider knots up to mutation. Secondly, a small family of pretzel knots, which we label exceptional, while they can for the most part be dealt with individually, has so far resisted all systematic efforts at ribbon classification; accordingly we restrict ourselves to the non-exceptional knots. Finally, if a pretzel knot is not prime, it on the one hand is a connected sum of two-bridge knots which obviously fiber [Gab86, Remarks 6.2], and thus itself fibers, while ribbonness of these was determined by Lisca [Lis07b].
@@
    $K=\pm P(1,1,1,1,-3,-3,-3)$ (note that this is ±$\pm 10_{75}$ in the Rolfsen table), or
@@
We remark that the knot ±$\pm 10_{75}$ is actually mutation-invariant, so the ordering of parameters in the case (1) is arbitrary. Also, knots with parameters in cases (1), (3) and (4) are never exceptional. It is also worth noting that in the literature pretzel knots are often classified as odd, i.e. defined by an odd number of odd parameters, or even otherwise. In this language, Theorem 1.1 implies that ±$\pm 10_{75}$ is in fact the unique fibered ribbon odd pretzel knot, while there are infinitely many even ones.
We want to stress that reordering of the parameters plays very differently with fiberedness and ribbonness, and we won’t attempt to pin down the different possibilites. To showcase the intricacy of the problem let us highlight that for example the pretzel knot $K=P(5,-5,7,-7,4)$ is ribbon and fibered; its mutant $K^{\prime}=P(7,5,-5,-7,4)$ is ribbon, but not fibered; furthermore $K^{\prime\prime}=P(7,-5,-7,5,4)$ is fibered but conjecturally not ribbon and finally $K^{\prime\prime\prime}=P(5,7,-5,-7,2)$ is neither fibered nor ribbon (conjecturally).
The strategy we follow to prove Theorem 1.1 is as follows: taking Gabai’s classification of fibered pretzel knots as a starting point, we determine which of them are potentially slice by obstructing their double covers from bounding rational homology balls using lattice embedding techniques derived from Donaldson’s Theorem A. The potentially slice candidates are then shown to be ribbon by explicitly describing the disks. It is worth noting that it is customary in the literature to exclude or to treat independently the case in which parameters can include ±$\pm 1$ or 00. In fact, many previous results, particularly regarding sliceness and ribonness properties, are either stated only for pretzel knots with no unitary parameters, or are treated separately [Lec15, Mil17, Bry17, Lon14]. This article will also treat the unitary and non-unitary cases separately. If 0 is a parameter, it can only appear once, and the knot is a connected sum, so as mentioned above falls into known classifications.
@@ 2. Pretzel knots: fiberedness and sliceness
Figure 1. The lefmotst figure, without the gray band, is the standard projection of the pretzel knot $P(-5,5,-3,3,7)$. The gray band represents the ribbon move that can always be performed when we have adjacent parameters of opposite sign and same absolute value. The other figures depict the result of performing the ribbon move.
@@
Parameters of absolute value one, referred to as unitary, can be permuted to any spot in a pretzel knot by flype moves. It follows that if $p_{i}=1$ and $p_{j}=-1$ in $K$, then $p_{i}$ and $p_{j}$ can be pairwise removed and $K=P(p_{1},\dots,\widehat{p_{i}},\dots,\widehat{p_{j}},\dots p_{n})$. For economy of notation, we will write $P([1^{\pm d}],p_{d+1},\dots,p_{n})$ to indicate the pretzel knot with the first $d$ parameters equal to ±$\pm 1$. We can, and from now on will, further assume that in $K$ we do not have indices $i$ and $j$ such that $p_{i}=\pm 1$ and $p_{j}=\mp 2$: by flyping we can simplify such a pretzel knot to $K=P(p_{1},\dots,\widehat{p_{i}},\dots,p^{\prime}_{j}=\pm 2,\dots,p_{n})$. The non-unitary parameters (i.e. with $|p_{i}|>1$) can be cyclically permuted and their order can be reversed without changing the knot type. Further permutations of the parameters might yield non isotopic pretzel knots. For example $P(3,5,7,2)$ is isotopic to $P(5,7,2,3)$ but not to $P(3,7,5,2)$. All changes of order in the parameters yield mutant pretzel knots, which share the double covers of $S^{3}$ branched over them [Bed84]. The mirror of the pretzel knot $K=P(p_{1},\dots,p_{n})$ is $\bar{K}=-P(p_{1},\dots,p_{n})=P(-p_{1},\dots,-p_{n})$.
A complete list of fibred pretzel links, with a description of their fibered surfaces, was established by Gabai in [Gab86]. In this article we only deal with knots, which simplifies Gabai’s result substantially. We summarize here the results relevant to us. For convenience, unless otherwise specified, we will order the parameters of knots of Types 2 and 3 such that the only even parameter is the last one. For these knots, Gabai’s classification uses an auxiliary pretzel link constructed as follows. For Type 2, record the sign of the odd non-unitary parameters as $a_{i}=\frac{p_{i}}{|p_{i}|}$ and to $K=P([1^{\pm d}],p_{d+1},\dots,p_{n-1},2m)$ associate the pretzel link $L^{\prime}=P(-2a_{d+1},-2a_{d+2},\dots,-2a_{n-1},2m)$. For Type 3 the construction of $L^{\prime}$ is very similar: this time we record the sign of all non-unitary parameters and $L^{\prime}$ becomes the link with only ±$\pm 2$ entries defined as $L^{\prime}=P(-2a_{d+1},-2a_{d+2},\dots,-2a_{n-1},\frac{2m}{|m|})$. Finally, we recall Gabai’s theorem, adapted to our setting and notation:
Theorem 2.1 (Gabai, Theorem 6.7 in [Gab86]).
@@ 1. Type 1:
    $K$ fibers if and only if each $p_{i}$ is ±$\pm 1$ or $\mp 3$ and some $p_{i}=\pm 1$. These knots are of the form ±$\pm P([1^{d}],-3,\dots,-3)$ with $d\geq 1$.
@@ 2. Type 3B:
      The numbers of positive and negative parameters are equal and $L^{\prime}\neq\pm P(2,-2,\dots,2,-2)$. Then $K$ fibers if and only if the auxiliary link $L^{\prime}=\pm P(2,-2,\dots,2,-2,-2)$.
@@ 3. Type 3C:
Because of the importance of the order of the parameters defining a pretzel knot, we want to stress that fiberedness is a subtle property when we consider re-ordering of parameters. For example $P(1,5,7,-5,-2)$ is a Type 2A fibered pretzel knot and all its mutants, that is, all the different Type 2 pretzel knots obtained by reordering the parameters, for example $P(1,-5,7,5,-2)$, are still fibered. On the other hand, if we consider a Type 2B fibered pretzel knot, such as $P(3,-7,5,-5,8)$, the order of the parameters is crucial: the mutant $P(3,5,-7,-5,8)$ is not fibered.
In this article, we aim to establish which fibered pretzel knots are ribbon. The general question of understanding which pretzel knots are slice and/or ribbon is very far from settled. Many of the known results use as a starting point the fact that the double cover of $S^{3}$ branched over a slice knot $K$, call it $Y_{K}$, bounds a rational homology ball. If one can show that $Y_{K}$ does not bound such a ball, then $K$ is not slice. When dealing with pretzel knots, this approach has the following interesting feature: if we start with a pretzel knot $K=P(p_{1},\dots,p_{n})$ and $Y_{K}$ does not bound a rational homology ball, we get for free that none of the mutants of $K$ is slice either. That is, we conclude that no matter what order we choose for the set of parameters $\{p_{1},\dots,p_{n}\}$, a pretzel knot with those parameters is not slice. On the other hand, if we show that $Y_{K}$ does bound a rational homology ball, and say we manage to exhibit a slice disk for $K=P(p_{1},\dots,p_{n})$, then all the other pretzel knots defined with the same set of parameters but a different order might not be slice but all the obstructions derived from the double cover will vanish. There is one exception to all this: 3-stranded pretzel knots are determined by their double covers, since cyclic permutation and reversal of order account for all possible orderings of three parameters.
The first systematic study of the slice ribbon conjecture for pretzel knots was carried out by Greene and Jabuka [GJ11] who dealt with 3-stranded Type 1 pretzel knots. They show that if no parameter is unitary, then such a pretzel is ribbon if and only if it is of the form $P(p,-p,q)$; if there are unitary parameters, then the pretzel knot is actually a 2-bridge knot and the slice ribbon conjecture has been proven for this class of knots in [Lis07a].
Chronologically, the next big step towards classifying slice pretzel knots was carried out in [Lec15], where it was shown that
@@
Here an exceptional pretzel knot is a Type 2 knot whose parameters consist of pairs of numbers of the same absolute value and opposite signs along with $(a,-a-1,-\frac{(a+1)^{2}}{2})$ with $a\equiv 1,97\mod(120)$ [Lec15, KLS22]. While these are conjecturally not even algebraically slice, they have resisted all attempts so far to establish this statement.
The case of 5-stranded Type 1 pretzel knots was addressed in [Bry17]. Particularly, Corollary 2 in that article tells us that in the absence of unitary parameters, if these pretzel knots are slice then the defining set of parameters is of the form $\{p_{0},p_{1},-p_{1},p_{2},-p_{2}\}$. At the time of writing, beyond 5-strands, the ribbon status of Type 1 pretzel knots was wide open.
@@
With the currently available techniques, dealing with mutation is difficult. One of the frequently used tools to distinguish mutants are twisted Alexander polynomials. These, while effective, involve intricate computations which do not play well with infinite families. In [HKL10] this technique was used to show that the 24 distinct oriented mutants of $P(3,7,9,11,15)$ are mutually different in the topological concordance group. Directly relevant to our exposition here is the result of Miller [Mil17, Theorem 1.4] who is able to show that infinitely many 4-stranded pretzels with ‘the right parameters but the wrong order’ are not slice.
The results in the literature pinning down the possible sets of parameters defining ribbon pretzel knots do say something about the order. In fact, [Lec15, Proposition 2.1] presents a very simple algorithm describing ribbon disks for pretzels with paired up parameters that satisfy an adjacency condition. For example, the algorithm applied to $P(3,-3,5,-5,7)$ yields a ribbon disk while it does not for the mutant $P(3,5,-3,-5,7)$. It was in fact shown in [HKL10] that this last pretzel knot is not slice. Conjecturally [Lec15, Conjecture 2.4], only the pretzel knots, without unitary parameters, for which the aforementioned algorithm yields a disk, are ribbon. In this article, we will call such pretzel knots detectably ribbon (Cf. ‘simple ribbon’ in [Bry17]).
We close this background section with a word about unitary parameters. In the following sections we will show that in their presence, ribbon pretzel knots do not exactly follow the pattern of ‘paired up parameters’ (see for example Propositions 4.1 and 4.7). Moreover, the bands describing the ribbon disks are not particularly simple in the standard pretzel projection of these knots (see Figures 6 and 13). These examples show that the “Pretzel Slice-Ribbon Conjecture” in [Bry17, Section 2] needs to be amended to exclude the case of unitary parameters.

### v2

@@ Abstract.
We classify fibered ribbon pretzel knots up to mutation. The classification is complete, except perhaps for members of Lecuona’s “exceptional” family of [pretzel]. The result is obtained by combining lattice embedding techniques with Gabai’s classification of fibered pretzel knots, and exhibiting ribbon disks, some of which lie outside of known patterns for standard pretzel projections.
@@ 1. Introduction
The intersection of fiberedness and ribbon/sliceness is of particular interest, in large part due to work of Casson and Gordon [casson1983loop], who observe that existence of a ribbon disc implies a surjection at the level of $\pi_{1}$ from the knot complement to the disk complement, which of course by definition implies sliceness. Labelling this algebraic condition homotopically ribbon, we have thus the following inclusions:
$\{\mathrm{ribbon}\}\subseteq\{\mathrm{homotopically\ ribbon}\}\subseteq\{\mathrm{slice}\}.$
Moreover, in the case of fibered knots, Casson and Gordon show that homotopical ribbonness admits a simple characterisation in terms of the fibration, namely that its monodromy extends over a handlebody. Investigation of the details of this result led Miller [Maggie, Remark 6.4] to study fibered ribbon 2-bridge knots, concluding that there are exactly 5 of these.
@@
To state our results, some conventions, standard when discussing pretzel knots, need to be established. Firstly, as we describe in detail in the next section, such knots can be encoded by an ordered tuple of integers. Mutations of the knot correspond to re-orderings of these parameters; as mutation preserves most known obstructions to sliceness, we will only consider knots up to mutation. Secondly, a small family of pretzel knots, which we label exceptional, while they can for the most part be dealt with individually, has so far resisted all systematic efforts at ribbon classification; accordingly we restrict ourselves to the non-exceptional knots. Finally, if a pretzel knot is not prime, it on the one hand is a connected sum of two-bridge knots which obviously fiber [Gabai, Remarks 6.2], and thus itself fibers, while ribbonness of these was determined by Lisca [lisca2007sums].
@@
    $K=\pm P(1,1,1,1,-3,-3,-3)$ (note that this is $\pm 10_{75}$ in the Rolfsen table), or
@@
We remark that the knot $\pm 10_{75}$ is actually mutation-invariant, so the ordering of parameters in the case (1) is arbitrary. Also, knots with parameters in cases (1), (3) and (4) are never exceptional. It is also worth noting that in the literature pretzel knots are often classified as odd, i.e. defined by an odd number of odd parameters, or even otherwise. In this language, Theorem 1.1 implies that $\pm 10_{75}$ is in fact the unique fibered ribbon odd pretzel knot, while there are infinitely many even ones.
We want to stress that reordering of the parameters plays very differently with fiberedness and ribbonness, and we won’t attempt to pin down the different possibilites. To showcase the intricacy of the problem let us highlight that for example the pretzel knot K=P(3,3,5,5,7,7,4)K=P(3,-3,5,-5,7,-7,4) is ribbon and fibered; its mutant $K^{\prime}=P(3,5,7,-7,-5,-3,4)$ is ribbon, but not fibered; furthermore $K^{\prime\prime}=P(3,-5,7,-3,5,-7,4)$ is fibered but, as Nathan Dunfield pointed out in private communication, can be shown using SnapPy’s implementation of Herald-Kirk-Livingston slice obstructions to not be ribbon [SnapPy], and finally the mutant $K^{\prime\prime\prime}=P(3,5,7,-3,-5,-7,4)$ is neither fibered nor ribbon.
The strategy we follow to prove Theorem 1.1 is as follows: taking Gabai’s classification of fibered pretzel knots as a starting point, we determine which of them are potentially slice by obstructing their double covers from bounding rational homology balls using lattice embedding techniques derived from Donaldson’s Theorem A. The potentially slice candidates are then shown to be ribbon by explicitly describing the disks. It is worth noting that it is customary in the literature to exclude or to treat independently the case in which parameters can include $\pm 1$ or 0. In fact, many previous results, particularly regarding sliceness and ribbonness properties, are either stated only for pretzel knots with no unitary parameters, or are treated separately [pretzel, miller, bryant, long]. This article will also treat the unitary and non-unitary cases separately. If 0 is a parameter, it can only appear once, and the knot is a connected sum, so as mentioned above falls into known classifications.
@@
Acknowledgements. We would like to thank Nathan Dunfield for helpful correspondence, and the anonymous referee for a remarkably careful reading of our article and their comments and suggestions.
@@ 2. Pretzel knots: fiberedness and sliceness
Figure 1. The leftmost figure, without the gray band, is the standard projection of the pretzel knot P(5,5,3,3,7)P(-5,5,-3,3,7). The gray band represents the ribbon move that can always be performed when we have adjacent parameters of opposite sign and same absolute value. The other figures depict the result of performing the ribbon move.
@@
Parameters of absolute value one, referred to as unitary, can be permuted to any spot in a pretzel knot by flype moves. It follows that if $p_{i}=1$ and $p_{j}=-1$ in $K$, then $p_{i}$ and $p_{j}$ can be pairwise removed and $K=P(p_{1},\dots,\widehat{p_{i}},\dots,\widehat{p_{j}},\dots p_{n})$. For economy of notation, we will write $P([1^{\pm d}],p_{d+1},\dots,p_{n})$ to indicate the pretzel knot with the first $d$ parameters equal to $\pm 1$. We can, and from now on will, further assume that in $K$ we do not have indices $i$ and $j$ such that $p_{i}=\pm 1$ and $p_{j}=\mp 2$: by flyping we can simplify such a pretzel knot to $K=P(p_{1},\dots,\widehat{p_{i}},\dots,p^{\prime}_{j}=\pm 2,\dots,p_{n})$. The non-unitary parameters (i.e. with $|p_{i}|>1$) can be cyclically permuted and their order can be reversed without changing the knot type. Further permutations of the parameters might yield non isotopic pretzel knots. For example $P(3,5,7,2)$ is isotopic to $P(5,7,2,3)$ but not to $P(3,7,5,2)$. All changes of order in the parameters yield mutant pretzel knots, which share the double covers of $S^{3}$ branched over them [bedient]. The mirror of the pretzel knot $K=P(p_{1},\dots,p_{n})$ is $\bar{K}=-P(p_{1},\dots,p_{n})=P(-p_{1},\dots,-p_{n})$.
A complete list of fibred pretzel links, with a description of their fibered surfaces, was established by Gabai in [Gabai]. In this article we only deal with knots, which simplifies Gabai’s result substantially. We summarize here the results relevant to us. For convenience, unless otherwise specified, we will order the parameters of knots of Types 2 and 3 such that the only even parameter is the last one. For these knots, Gabai’s classification uses an auxiliary pretzel link constructed as follows. For Type 2, record the sign of the odd non-unitary parameters as $a_{i}=\frac{p_{i}}{|p_{i}|}$ and to $K=P([1^{\pm d}],p_{d+1},\dots,p_{n-1},2m)$ associate the pretzel link $L^{\prime}=P(-2a_{d+1},-2a_{d+2},\dots,-2a_{n-1},2m)$. For Type 3 the construction of $L^{\prime}$ is very similar: this time we record the sign of all non-unitary parameters and $L^{\prime}$ becomes the link with only $\pm 2$ entries defined as $L^{\prime}=P(-2a_{d+1},-2a_{d+2},\dots,-2a_{n-1},\frac{2m}{|m|})$. Finally, we recall Gabai’s theorem, adapted to our setting and notation:
Theorem 2.1 (Gabai, Theorem 6.7 in [Gabai]).
@@ 1. Type 1:
    $K$ fibers if and only if each $p_{i}$ is $\pm 1$ or $\mp 3$ and some $p_{i}=\pm 1$. These knots are of the form $\pm P([1^{d}],-3,\dots,-3)$ with $d\geq 1$.
@@ 2. Type 3B:
      The numbers of positive and negative parameters are equal and $L^{\prime}\neq\pm P(2,-2,\dots,2,-2)$. Then $K$ fibers if and only if the auxiliary link is $L^{\prime}=\pm P(2,-2,\dots,2,-2,-2)$.
@@ 3. Type 3C:
Because of the importance of the order of the parameters defining a pretzel knot, we want to stress that fiberedness is a subtle property when we consider re-ordering of parameters. For example P(1,5,7,5,2)P(1,5,7,-5,-2) is a Type 2A fibered pretzel knot and all its mutants, that is, all the different Type 2 pretzel knots obtained by reordering the parameters, for example P(1,5,7,5,2)P(1,-5,7,5,-2), are still fibered. On the other hand, if we consider a Type 2B fibered pretzel knot, such as P(3,7,5,5,8)P(3,-7,5,-5,8), the order of the parameters is crucial: the mutant P(3,5,7,5,8)P(3,5,-7,-5,8) is not fibered.
In this article, we aim to establish which fibered pretzel knots are ribbon. The general question of understanding which pretzel knots are slice and/or ribbon is very far from settled. Many of the known results use as a starting point the fact that the double cover of $S^{3}$ branched over a slice knot $K$, call it $Y_{K}$, bounds a rational homology ball. If one can show that $Y_{K}$ does not bound such a ball, then $K$ is not slice. When dealing with pretzel knots, this approach has the following interesting feature: if we start with a pretzel knot $K=P(p_{1},\dots,p_{n})$ and $Y_{K}$ does not bound a rational homology ball, we get for free that none of the mutants of $K$ is slice either. That is, we conclude that no matter what order we choose for the set of parameters $\{p_{1},\dots,p_{n}\}$, a pretzel knot with those parameters is not slice. On the other hand, if we show that $Y_{K}$ does bound a rational homology ball, and say we manage to exhibit a slice disk for $K=P(p_{1},\dots,p_{n})$, then all the other pretzel knots defined with the same set of parameters but a different order might not be slice but this particular obstruction derived from the double cover will vanish. There is one exception to all this: 3-stranded pretzel knots are determined by their double covers, since cyclic permutation and reversal of order account for all possible orderings of three parameters.
The first systematic study of the slice ribbon conjecture for pretzel knots was carried out by Greene and Jabuka [GJ] who dealt with 3-stranded Type 1 pretzel knots. They show that if no parameter is unitary, then such a pretzel is ribbon if and only if it is of the form P(p,p,q)P(p,-p,q); if there are unitary parameters, then the pretzel knot is actually a 2-bridge knot and the slice ribbon conjecture has been proven for this class of knots in [lisca].
Chronologically, the next big step towards classifying slice pretzel knots was carried out in [pretzel], where it was shown that
@@
Here an exceptional pretzel knot is a Type 2 knot whose parameters consist of pairs of numbers of the same absolute value and opposite signs along with $(a,-a-2,-\frac{(a+1)^{2}}{2})$ with $a\equiv 1,97\mod(120)$ [pretzel, KimLeeSong]. While these are conjecturally not even algebraically slice, and on a case by case basis it is easy to establish that they are not, they have resisted all attempts so far to establish this statement for all parameters $a$.
The case of 5-stranded Type 1 pretzel knots was addressed in [bryant]. Particularly, Corollary 2 in that article tells us that in the absence of unitary parameters, if these pretzel knots are slice then the defining set of parameters is of the form $\{p_{0},p_{1},-p_{1},p_{2},-p_{2}\}$. At the time of writing, beyond 5-strands, the ribbon status of Type 1 pretzel knots was wide open.
@@
With the currently available techniques, dealing with mutation is difficult. One of the frequently used tools to distinguish mutants are twisted Alexander polynomials. These, while effective, involve intricate computations which do not play well with infinite families. In [HeraldKirkLiv] this technique was used to show that the 24 distinct oriented mutants of $P(3,7,9,11,15)$ are mutually different in the topological concordance group. Directly relevant to our exposition here is the result of Miller [miller, Theorem 1.4] who is able to show that infinitely many 4-stranded pretzels with ‘the right parameters but the wrong order’ are not slice.
The results in the literature pinning down the possible sets of parameters defining ribbon pretzel knots do say something about the order. In fact, [pretzel, Proposition 2.1] presents a very simple algorithm describing ribbon disks for pretzels with paired up parameters that satisfy an adjacency condition. For example, the algorithm applied to P(3,3,5,5,7)P(3,-3,5,-5,7) yields a ribbon disk while it does not for the mutant P(3,5,3,5,7)P(3,5,-3,-5,7). It was in fact shown in [HeraldKirkLiv] that this last pretzel knot is not slice. Conjecturally [pretzel, Conjecture 2.4], only the pretzel knots, without unitary parameters, for which the aforementioned algorithm yields a disk, are ribbon. In this article, we will call such pretzel knots detectably ribbon (Cf. ‘simple ribbon’ in [bryant]).
We close this background section with a word about unitary parameters. In the following sections we will show that in their presence, ribbon pretzel knots do not exactly follow the pattern of ‘paired up parameters’ (see for example Propositions 4.1 and 4.7). Moreover, the bands describing the ribbon disks are not particularly simple in the standard pretzel projection of these knots (see Figures 2 and 9). These examples show that the “Pretzel Slice-Ribbon Conjecture” in [bryant, Section 2] needs to be amended to exclude the case of unitary parameters.